OpenStudy (surana):

The graph of y < -x2 +4x - 4 contains which of the following points?

OpenStudy (surana):

And the possible answers are... A: 0,0 B: -3,2 C: 1,5 D: None of the above

OpenStudy (ashes.boom):

Never mind i have to go sorry

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

we have to test them out

OpenStudy (surana):

Hello, good to see you again. I was wondering if you could help here?

OpenStudy (misty1212):

sure lets try \((0,0)\) which i know does not work check if \[0<0^2+4\times 0-4\] is true, but it is not since \[0<-4\] is FALSE

OpenStudy (surana):

I'm guessing we have to plug in the numbers into the x and y in the order of x and y, right?

OpenStudy (surana):

But I do agree with 0,0 being wrong.

OpenStudy (misty1212):

lets try \((-3,2)\)

OpenStudy (surana):

So I put -3,2 into the other segments, right? So -3 into x^2 and 2 into x?

OpenStudy (misty1212):

\[2<(-3)^2+4\times (-3)-4\]\ or \[2<9-12-4\\ 2<-7\] also false

OpenStudy (surana):

So if I plug those in, they'll turn up wrong, so A and B are wrong ?

OpenStudy (misty1212):

yes

OpenStudy (surana):

So that leaves us with 1,5 or none of the above.

OpenStudy (misty1212):

i think none of them work

OpenStudy (misty1212):

\[5<1^2+4\times 1-4\\ 5<1\] also false

OpenStudy (surana):

So the answer would be none of the above, then?

OpenStudy (misty1212):

i think so yes

OpenStudy (surana):

Thank you, Misty.

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]